The product of two numbers is 4 square root of 3. Find the numbers so that the

sum š¯‘† of the square of one and the cube of the other is as small as

possible


1
Expert's answer
2021-06-15T09:19:33-0400

let x and y be two numbers.

The product of two numbers,

"xy=4\sqrt3\\
x=\frac{4\sqrt3}{y}"-----------(1)

sum š¯‘† of the square of one and the cube of the other is as small as

possible

"S=x^2+y^3\\
=(\frac{4\sqrt3}{y})^2+y^3\\
=\frac{48}{y^2}+y^3"

Here, S is the objective function.

"S'=\frac{-48}{y}+3y^2\\
Now, S'=0\\
\frac{-48}{y}+3y^2=0\\
y=2.5
\implies x=\frac{4\sqrt3}{2.5}=2.7"

thus, x=2.7 and y=2.5